OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to the line y= 3/5x - 1 and passes through the point (-9, 4). I have chosen: A in this case. A: y= 5/3x +19 B: y= 3/5x + 47/5 C: y= -5/3x - 11 D: y= 5/3x + 19 I am just trying to get the best grade possible and want to know if I am wrong...I only have one more after this and I should be fine ;-; Halps pls

OpenStudy (anonymous):

It is C, because it is perpendicular. It is opposite reciprocal.you did the reciprocal well, but didnt do the opposite part.

OpenStudy (anonymous):

Thanks :c I am going to have to look over the rest of this now because I am scared for the actual quiz I need to take soon.. <3

OpenStudy (anonymous):

:) Good Luck with that!!!!!

OpenStudy (anonymous):

If you need help, just ask me

OpenStudy (anonymous):

I do have a last question in my text book for my studies? Could you tell me if I am right ;-; Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) A: y = -2/3x + 3 B: y = -2/3x + 5/3 C: y = 2x - 21 D: y = -2/3x - 9 I got B when I worked it out :3

OpenStudy (anonymous):

yasssss great!!!

OpenStudy (anonymous):

Omg yaaay, I got this cx Thank you c: